# Flexible Tails for Normalizing Flows

### v1

@@ D.2 Sampling Student’s T
1:Input: degrees of freedom $\nu$, threshold $\epsilon\geq 0$.
2:Sample $g\sim\textrm{Gamma}(\nu/2,1)$.
3:Let $g^{\prime}=\max(g,\epsilon)$.
4:Sample $z\sim\mathcal{N}(0,1)$.
5:Return $z\sqrt{\frac{\nu}{2g^{\prime}}}$.
@@
D.3 Inverse Transformation for Small Inputs
@@
Appendix E Variational Inference Diagnostics
We measure the quality of a density $q_{x}(x)$ produced by VI using diagnostics based on importance sampling. This involves calculating $w_{i}=p(x_{i})/q_{x}(x_{i})$ where $p(x)$ is the target density and $x_{i}\sim q_{x}(x)$ (independently) for $i=1,2,\ldots,n$. In our examples we use $n=10,000$.
@@ Appendix E Variational Inference Diagnostics
Our second diagnostic is from Yao et al. (2018), who fit a GPD to the $w_{i}$s and return the estimated shape parameter $\hat{k}$. Lower $\hat{k}$ values indicate a better approximation of $p(x)$ by $q_{x}(x)$. Yao et al. (2018) argue that $\hat{k}>0.7$ corresponds to a very poor approximation, and $\hat{k}<0.5$ indicates a good fit.
Appendix F Further Discussion of COMET Flows
Performance
Where we can implement COMET flows in our empirical comparisons, they are competitive with the best performing methods—see Table 2. However our theory in Appendix B.2 shows they can’t capture all heavy tails, in one case producing log-normal tails rather than GPD tails. One possible reason for good empirical behaviour is that log-normal and GPD distributions have similar sub-asymptotic properties (Nair et al., 2022).
Limitations
In our density estimation example on synthetic data (Section 4.2), we found that COMET flows were too slow to implement for $d=50$. It’s unclear if the issue is fundamental to the method or due to its implementation. For instance, the existing implementation requires repeatedly moving data between the CPU and GPU, so a more efficient implementation might be more competitive.
We also found some bugs in the existing implementation of COMET flows, which are fixed in the code for our paper.
COMET flows involve an initial stage of KDE estimation of marginal distributions. This is not usually possible for VI, so the method is not generally applicable there.

### v2

@@ G.4 Sampling Student’s T
0:  Input: degrees of freedom $\nu$, threshold $\epsilon\geq 0$.
1:  Sample $g\sim\text{Gamma}(\nu/2,1)$.
2:  Let $g^{\prime}=\max(g,\epsilon)$.
3:  Sample $z\sim\mathcal{N}(0,1)$.
4:  Return $z\sqrt{\frac{\nu}{2g^{\prime}}}$.
@@
G.5 Inverse Transformation for Small Inputs
@@
G.6 Optimisation
All of the experiments use the Adam optimiser. Table 5 gives learning rate and batch size values.
Most experiments use pytorch defaults for other tuning choices. One exception is density estimation for CLIMDEX, where we follow the optimisation setup from Laszkiewicz et al. (2022). Another is the $\nu=0.5$ variational inference example, where we clip the gradient norm to 5, which was beneficial to the stability of mTAF and TTFfix.
Table 5: Optimisation Hyperparameters.
Experiment Group Learning Rate Batch Size
Synthetic Density Estimation (Section 4.1) 5e-3 None (full pass)
Real Data (Section 4.2) 5e-4 512
Variational Inference (Section 4.3) 1e-3 100
Appendix H Variational Inference Diagnostics
We measure the quality of a density $q_{x}(x)$ produced by VI using two diagnostics based on importance sampling. This involves calculating $w_{i}=p(x_{i})/q_{x}(x_{i})$ where $p(x)$ is the target density and $x_{i}\sim q_{x}(x)$ (independently) for $i=1,2,\ldots,n$. In our examples we use $n=10,000$.
@@ Appendix H Variational Inference Diagnostics
Our second diagnostic is from Yao et al. (2018), who fit a GPD to the $w_{i}$s and return the estimated shape parameter $\hat{k}$. Lower $\hat{k}$ values indicate a better approximation of $p(x)$ by $q_{x}(x)$. Yao et al. (2018) argue that $\hat{k}<0.7$ indicates a useful variational approximation, so we highlight this threshold in our tables.
Appendix I Data
Table 6 summarises the real data sets used in our experiments.
Table 6: Real data sets information.
Name Dimension Average $\nu$ Topic Source
Insurance 2 2.17 Medical claims Liang et al. (2022)
Fama 5 5 2.36 Daily returns of 5 major indices Liang et al. (2022)
S&P 500 300 4.78 Daily returns of the 300 most traded US stocks Novel to our paper
CLIMDEX 412 4.24 High dimensional meteorological data Laszkiewicz et al. (2022)
All datasets are pre-processed before density estimation is performed. To do so, data is normalized to zero mean and unit variance using the estimated mean and variance from the training and validation sets, as is standard practice in NF density estimation (Papamakarios et al., 2017; Durkan et al., 2019).
I.1 S&P 500
Here we describe the S&P 500 dataset which we introduce. These financial returns are an example of moderately high dimensional multivariate data with extreme values. We take the closing prices of the top 300 most traded S&P 500 stocks, and convert them in standard fashion to log returns i.e. the log returns are $x_{j}=\log(\frac{s_{j+1}}{s_{j}})$ where $s_{j}$ is stock closing price on day $j$. We use data covering the time period 2010-01-04 to 2022-10-27, corresponding to 3227 days in total. In this example we concentrate on the tails of the data, rather than time series structure. As such, we treat each day of log returns as an independent observation in $\mathbb{R}^{d}$.
The test set is comprised of observations after 2017-09-14, with train and validation sampled uniformly from the period up to and including this date. This corresponds to 1292 training, 645 validation and 1290 test observations respectively.
Appendix J Additional Results for Density Estimation with Synthetic Data
J.1 Other Experimental Settings
Table 7 expands on the results from Table 1 from the main paper, by including more values of $d$ and $\nu$.
Dashes in the table indicate potential unstable optimisation: at least one repeat had a very large final loss (above 1e5). In these cases, even excluding losses this large resulted in significantly worse mean losses than other methods. These results confirm that all the methods not specifically designed to permit GPD tails become increasingly unstable as the data becomes more heavy tailed.
The table includes results for the TTF_tBase architecture. As described in Section 4.1, this is similar to TTF, but uses a Student’s T base distribution with trainable degrees of freedom. TTF_tBase always does worse than TTF and TTFfix. For the largest $\nu$ values – $2$ and $30$ – it is the worst of all methods. It also has the largest standard error values, suggesting that optimisation may be more difficult once these two methods are combined and there are two sets of tail parameters to tune.
Table 7: Density estimation results on synthetic example. Each entry is a mean value of negative test log likelihood per dimension across 10 repeated experiments, with the standard error in brackets. (Dividing by dimension acts to normalises the values, allowing for easier comparison across dimensions.) Bold indicates methods whose mean log likelihood differs from the best mean by less than 2 standard errors (of the best mean). A dash indicates potential unstable optimisation (at least one repeat had a final loss above 1e5). No entry indicates that the experiment was not ran.
$d$ Flow $\nu=0.5$ $\nu=1$ $\nu=2$ $\nu=30$
5 normal - - 2.01 (0.07) 1.46 (0.00)
m_normal - 403.24 (239.15) 1.94 (0.02) 1.46 (0.00)
g_normal - - 1.93 (0.01) 1.46 (0.00)
gTAF 6.42 (0.27) 2.49 (0.02) 1.90 (0.01) 1.46 (0.00)
TTF 3.33 (0.01) 2.34 (0.01) 1.89 (0.01) 1.47 (0.00)
mTAF 4.08 (0.03) 2.49 (0.02) 1.92 (0.01) 1.46 (0.00)
TTFfix 3.33 (0.01) 2.35 (0.01) 1.89 (0.01) 1.47 (0.00)
COMET 3.42 (0.01) 2.35 (0.01) 1.89 (0.00) 1.46 (0.00)
TTF_tBase 3.39 (0.02) 2.51 (0.04) 2.04 (0.02) 1.65 (0.03)
10 normal - - 2.00 (0.02) 1.46 (0.00)
m_normal - - 2.04 (0.07) 1.46 (0.00)
g_normal - - 1.98 (0.02) 1.46 (0.00)
gTAF 7.13 (0.31) 2.63 (0.02) 1.95 (0.00) 1.47 (0.00)
TTF 3.55 (0.01) 2.47 (0.00) 1.93 (0.00) 1.47 (0.00)
mTAF 4.48 (0.04) 2.63 (0.01) 1.95 (0.00) 1.46 (0.00)
TTFfix 3.54 (0.01) 2.46 (0.00) 1.93 (0.00) 1.47 (0.00)
COMET 3.63 (0.01) 2.46 (0.00) 1.93 (0.00) 1.47 (0.00)
TTF_tBase 3.67 (0.01) 2.63 (0.01) 2.08 (0.01) 1.62 (0.01)
50 normal - - 2.02 (0.01) 1.47 (0.00)
m_normal - - 2.02 (0.00) 1.47 (0.00)
g_normal - - 2.01 (0.00) 1.47 (0.00)
gTAF 7.49 (0.38) 2.65 (0.01) 1.99 (0.00) 1.47 (0.00)
TTF 3.68 (0.00) 2.54 (0.00) 1.98 (0.00) 1.47 (0.00)
mTAF 5.22 (0.04) 2.62 (0.01) 1.98 (0.00) 1.47 (0.00)
TTFfix 3.68 (0.00) 2.54 (0.00) 1.98 (0.00) 1.47 (0.00)
COMET 3.74 (0.00) 2.55 (0.00) 1.97 (0.00) 1.47 (0.00)
TTF_tBase 4.17 (0.01) 2.84 (0.04) 2.35 (0.04) 1.82 (0.06)
J.2 Learned Tail Parameters
Figure 2 illustrates tail parameters learned by TTF corresponding to the $d=5$ experiment of Table 7. These results show that the learned parameters don’t exactly match the TTFfix values chosen to match the known tail shapes. Our theory – Theorem 12 statement 2 – shows that the two should be equal for a good tail fit if our transformation were applied to a $\mathcal{N}(0,1)$ random variable. However Theorem 12 statement 1 shows that changing the input random variable – e.g. to $\mathcal{N}(\mu,\sigma^{2})$ – can produce a resulting distribution with a different tail shape. So the lack of a match in practice suggests the learned TTF parameters adjust slightly from the TTFfix values to compensate for the effect of the other normalising flow layers.
Adjusting in this way could be an advantage in learning the tail parameters (as in our TTF method) rather than fixing them in advance (as in our TTFfix method). However in this case any such advantage is small, as the overall results are very similar in Table 7. Also, as we have seen in other experiments, TTFfix is often competitive in practice, and can be easier to train.
Figure 2: Box plots of TTF final tail shape parameters $\lambda$ from repeated optimiser runs. The dashed horizontal lines show the TTFfix values. The $y$-axis shows $1/\lambda$, as then the TTFfix values directly match the target Student’s $T$ degrees of freedom.